OpenStudy (anonymous):

Two cars leave the same location at the same time and travel along the same route to the same destination, which is 120 miles away. If car A arrived at the destination in 3 hours, and car B arrived 2 hours after car A, what was the average difference in speed, measured in miles per hour, of car A and car B?

OpenStudy (anonymous):

speedA = 120/3 speeB = 120/5 find speedA - speedB

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

you made it so simple, thank you so much!

OpenStudy (anonymous):

Awsuummm, correct You're welcome